Determine the truth value of each of the quantified statements below if the domain consists of R.
(a)∃x(x^5 = -1)
(b)∃x(x^6< x^4)
(c)∀x((-x)4=x^4)
(d)∀x(2x > x)
1
Expert's answer
2020-12-07T05:54:54-0500

(a) True. There is a x in R such that "x^5=-1" , it's x=-1.

(b) True. "x^6<x^4,x^4(x^2-1)<0,x^4<0" has no solutions so "x^2-1<0,|x|<1,-1<x<1". So yes, there is x in R such that "x^6<x^4" .

(c) True. No matter what x is, "(-x)^4=((-1)\cdot x)^4=(-1)^4\cdot x^4=1\cdot x^4=x^4" .

(d) False. "2x>x, x>0" . So this statement is true only for positive numbers, not for all x in R